OpenStudy (anonymous):

Check whether solutions of the following ivp exist: \[y''+e ^{x}y=x ^{2}\] --- y(0)=1, y'(0)=0

OpenStudy (amistre64):

is there a certain chk to perform, or is the intent to solve the diffyQ and check?

OpenStudy (anonymous):

I'm not sure. The problem says no to solve the equation. Do you plug in the initial conditions? But if so you don't have a y''.

OpenStudy (amistre64):

then theres some thrm or another that relates to this from your related text

OpenStudy (amistre64):

a wronskian is about all i can think of at the moment, but im sure that wouldnt be applicable

OpenStudy (anonymous):

Is there someway to check without actually solving the problem?

OpenStudy (amistre64):

im sure there is, which would relate to some thrm or another from the text that goes along with the material. Im just not too sure what that thrm might be http://www.wolframalpha.com/input/?i=y%E2%80%B2%E2%80%B2%2Be%5Ex+y%3Dx%5E2

OpenStudy (amistre64):

http://math.bard.edu/~mbelk/math308/solutions1.pdf this seems to ahve a few thrms

OpenStudy (amistre64):

seems that taking the df/dy of the y'' rewrite might be helpful

OpenStudy (anonymous):

What do you mean? y'' doesn't have a variable in front

OpenStudy (amistre64):

y′′ = x^2 - e^x y df/dy = -e^x

OpenStudy (amistre64):

since this is continuous at all y values, and all x values ...

OpenStudy (anonymous):

I believe that would only apply if the equation was non linear

OpenStudy (amistre64):

without any thrms from your text that relates to this, it was a shot in the dark :)

OpenStudy (amistre64):

good luck tho

OpenStudy (anonymous):

Thanks, I'll see if I can find something